OpenStudy (blackops2_143):

Solve for x. −2/3(3x−4)+3x=5/6 ​​ ​ x=−1/96 ​ ​ ​ x=−1/16 ​ ​ ​ x=2/16 ​ x=2/96

alones (alones):

OKay Simplify both sides of the equation \[\frac{ -2 }{ 3}(3x−4)+3x=\frac{ 5 }{ 6 }\]

alones (alones):

So, after disturbuting and combining like terms you should get this @blackops2_143 \[x+\frac{ 8 }{ 3 }=\frac{ 5 }{ 6 }\]

OpenStudy (blackops2_143):

okay thanks

alones (alones):

and finally to get to your answer you'll have to subtract \(\frac{ 8 }{ 3 }\)from both sides

OpenStudy (blackops2_143):

okay thank you ^.^

alones (alones):

\[x+\frac{ 8 }{ 3 }-\frac{ 8 }{ 3 }=\frac{ 5 }{ 6}-\frac{ 8 }{ 3 }\]

alones (alones):

I'm not done yet lolol anyways, subtract it, and tell me whatcha get, and i'll compare mine answer to yours (;

OpenStudy (rrhtgre):

-11/6 is the answer

alones (alones):

@rrhtgre you can't just give a direct answer, noob

OpenStudy (rrhtgre):

why not

OpenStudy (blackops2_143):

I was just about to type it but that's what I got

OpenStudy (blackops2_143):

thank you for the help!

alones (alones):

Yw hun (;